# CultureLLM: Incorporating Cultural Differences into Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are reported to be partial to certain cultures owing to the training data dominance from the English corpora. Since multilingual cultural data are often expensive to collect, existing efforts handle this by prompt engineering or culture-specific pre-training. However, they might overlook the knowledge deficiency of low-resource culture and require extensive computing resources. In this paper, we propose CultureLLM, a cost-effective solution to incorporate cultural differences into LLMs. CultureLLM adopts World Value Survey (WVS) as seed data and generates semantically equivalent training data via the proposed semantic data augmentation. Using only $50$ seed samples from WVS with augmented data, we fine-tune culture-specific LLMs and one unified model (CultureLLM-One) for $9$ cultures covering rich and low-resource languages. Extensive experiments on $60$ culture-related datasets demonstrate that CultureLLM significantly outperforms various counterparts such as GPT-3.5 (by $8.1$%) and Gemini Pro (by $9.5$%) with comparable performance to GPT-4 or even better. Our human study shows that the generated samples are semantically equivalent to the original samples, providing an effective solution for LLMs augmentation.
@@ 1 Introduction
Culture is a complex construct that encodes various identities, including but not limited to, language, nationality, region, religion, and gender identity. Cultural bias is widely present around the world, which refers to the tendency to favor specific cultural perspectives, values, and norms that lead to subjective opinions and can offend people of other cultures. For example, according to the World Value Survey (Survey, 2022), Arabic culture believes that men are better political leaders than women, while people in the United States disagree. [^3] We respect all opinions in different cultures. As large language models (LLMs) (OpenAI, 2023b; Google, 2023) become prevalent, they are reported to suffer cultural bias and are specifically partial to Western culture, as English corpora dominate training data (Liu et al., 2023b; Cao et al., 2023; Masoud et al., 2023; Naous et al., 2023; Wang et al., 2023d; Johnson et al., 2022). Low-resource cultures exist widely due to the lack of training data from other cultures. LLMs’ cultural bias presents a major bottleneck in human-AI collaboration and significantly hinders AI democracy.
Figure 1: Overview of CultureLLM. CultureLLM consists of three steps: sampling, semantic data augmentation, and fine-tuning. Both culture-specific and unified CultureLLM can be fine-tuned.
@@
Tackling cultural bias requires an LLM to embrace cultural differences (Hofstede, 1984). Kovač et al. (2023) and Wang et al. (2023d) thought LLMs have enough knowledge of all cultures and devised prompt engineering technologies to induce LLMs to exhibit specific cultural perspectives. However, they are not effective, especially in low-resource cultures with limited data. Another line of work pre-trained culturally-aware LLMs and then fine-tuned on specific datasets (Chan et al., 2023; Nguyen et al., 2023b; Pipatanakul et al., 2023; Abbasi et al., 2023; Lin and Chen, 2023). They require the collection of large-scale pre-training and fine-tuning datasets and extensive computing resources, thus are not affordable to ordinary researchers and cannot handle low-resource culture. To date, training culturally-aware LLMs at affordable costs remains a challenge.
@@
In this paper, we propose CultureLLM, a cost-effective [^4] Fine-tuning a CultureLLM only costs \$6 via OpenAI API. solution to incorporate cultural differences into LLMs. To be specific, we focus on cultural values in this work. As shown in Figure 1, CultureLLM consists of three steps: sampling, semantic data augmentation, and fine-tuning. Inspired by Attitude-Behavior Consistency theory (Fazio and Zanna, 1981) which emphasizes that people’s opinion is consistent with their behaviors, we use the World Values Survey (WVS) (Survey, 2022) as seed data. Then, we devise a semantic data augmentation approach to generate semantically equivalent samples. The aim is to generate semantic equivalent inputs, thus we can get the ground-truth from seed data directly. Finally, CultureLLM is obtained by fine-tuning on both the seed and the generated data. WVS is a public opinion poll that contains people’s opinions on cultural topics from different countries. To be specific, we select $50$ seed samples from WVS, covering $7$ topics: “social values", “migration", “security", “science and technology", “religious values", “ethical values and norms", and “political interest and political participation". Using these generated samples and answers from people in different cultures, we fine-tune specific and unified LLMs: specific LLMs are tailored for each culture such as CultureLLM-Ar for Arabic and CultureLLM-Tr for Turkish; unified LLMs (CultureLLM-One) are one LLM that fits all cultures. [^5] The bound of culture is unclear, and we use the main spoken language to distinguish cultures in this work (Delanoy, 2020).
@@
We build $9$ specific CultureLLM and a CultureLLM-One covering both high- and low-resource cultures: Arabic culture, Bengali culture, Chinese culture, English culture, German culture, Korean culture, Portuguese culture, Spanish culture, and Turkish culture. Then we evaluated them on $8$ culture-related downstream tasks: offensive language detection, hate speech detection, stance detection, toxicity detection, threat detection, bias detection, abusive detection, spam detection, and an open-ended generative task. We have $60$ test sets, which contain $68,672$ test samples in total. Experiments show that CultureLLM fine-tuned on GPT-3.5 significantly outperforms GPT-3.5 by $8.1$% and outperforms Gemini pro (Google, 2023) by $9.5$% on average F1 score, achieving comparable or even better performance with GPT-4. Our human study of $50$ people demonstrates that the augmentation method can generate semantically equivalent samples. We further interpret the rationale behind its effectiveness by exploring the fine-tuning data size and case studies. Finally, results on Big-Bench Hard (Suzgun et al., 2022) and GSM8K (Cobbe et al., 2021) indicate that CultureLLM is resistant to catastrophic forgetting. CultureLLM also supports open-source fine-tuning.
@@
    We conducted extensive experiments across a wide range of cultures and LLMs, showing that LLMs performs consistently well in all downstream tasks.
@@ 2.1 Cultural Problem and Solution in LLMs
The ideal solution is to enhance the cultural awareness of LLMs. There are mainly two types of approach: prompt engineering and pre-training. Kovač et al. (2023); Wang et al. (2023d) thought LLMs as superpositions of cultural perspectives, which can be prompted to targeted cultural perspectives. while Rao et al. (2023) encoded cultural values in the prompts. Although PE is cheap, its effectiveness is challenged, especially in low-resource cultures where LLMs lack such cultural knowledge due to lack of representation in pre-training data. Another line of research is pre-training and fine-tuning (Chan et al., 2023; Nguyen et al., 2023b; Pipatanakul et al., 2023; Abbasi et al., 2023; Lin and Chen, 2023) that trains culturally-aware LLMs for different cultures by collecting large-scale pre-training datasets and then performed fine-tuning for better alignment. While they achieved great performance, this approach is too expensive and time-consuming, thus it is difficult to apply to more cultures and countries. They still suffer from a low-resource culture problem where the pre-training data are difficult to collect. MaLA-500 (Lin et al., 2024) trained a new LLM on Llama 2 to cover $534$ languages, which is resource intensive.
@@ 2.2 Data Augmentation for LLMs
Human-annotated data are high-quality but expensive. Due to the strong generation ability of LLMs, many works focused on data augmentation leveraging LLMs.  Yu et al. (2023); Liu et al. (2023a) used LLMs to augment the math data and then fine-tuned with those data. Li et al. (2023) synthesized data with two devised modules: self-augmentation and self-curation. Chen et al. (2024) introduced a self-play mechanism, where LLM generates its own training data from its previous iterations, refining its policy by discerning these self-generated responses from those obtained from human-annotated data. There are also other uses for synthetic data, such as knowledge distillation (Wang et al., 2023c) and improving text embedding tasks (Wang et al., 2023a). Our data augmentation approach also adopts LLMs for data generation, but we add controllable modules such as template editing, synonym replacement, and semantic filter to ensure the diversity and semantic equivalence of the generated samples. Our approach can also be used as a general augmentation method in other applications.
2.3 Culture-related Research and Value Alignment
Efforts in cultural datasets (Nguyen et al., 2023a; Fung et al., 2022) focus on cultural common sense and norms, respectively. However, they generate data from only the English or Chinese corpus and thus may contain cultural bias toward other cultures. In contrast, World Values Survey (WVS) (Survey, 2022) is a large-scale pool that contains answers from vast people of different cultures, thus providing more objective cultural values from specific cultures.
This work is also related to value alignment (Ji et al., 2023; Shen et al., 2023; Yao et al., 2023), which focuses on aligning the values of LLMs with human’s by designing algorithms for value measurement and behavior alignment. In contrast, this work primarily emphasizes value understanding with the potential to be extended for value alignment. For instance, semantic augmentation can be used to generate training data for alignment-related tasks.
@@ 3 CultureLLM
In this section, we introduce how to build CultureLLM by fine-tuning existing LLMs via semantic data augmentation.
@@ 3.1 Overview
Cultural differences are prevalent in various cultures and backgrounds, leading to an impact on outcomes in downstream applications such as hate speech and biased language. To address the gap between low-source cultural data collection and its wide applications, we design CultureLLM by fine-tuning an LLM on data generated by our novel semantic data augmentation approach. Figure 1 presents an overview of CultureLLM, where the first step is to sample a subset of data from an existing World Value Survey (WVS) (Survey, 2022) that represents different opinions (answers) towards the same value questions given by native users. The adoption of WVS is inspired by Attitude-Behavior Consistency theory (Fazio and Zanna, 1981), which emphasizes the strong relationship between attitude and behavior. Therefore, WVS serves as an ideal seed for data augmentation. [^6] WVS is only one feasible option for seed and other surveys can also be used. But, as we discussed, the cultural survey data is extremely rare and WVS could be the most comprehensive one. After sampling, the second step is to generate augmented data using our proposed semantic augmentation approach (Section 3.3) and then fine-tune a CultureLLM for each specific culture such as CultureLLM-Ar for Arabic culture and CultureLLM-Tr for Turkish culture.
Generally speaking, we use $\mathcal{D}_{d}=\{(x_{j},y_{j}^{d})\}_{j=1}^{n}$ to denote the seed and $\mathcal{D}^{\prime}_{d}=g(\mathcal{D}_{d})=\{(x^{\prime}_{j},y_{j}^{d})\}_{j=1}^{n^{\prime}}$ as the augmented data with $g(\cdot)$ the augmentation algorithm. Note that the question $x$ here is the same in all cultures in WVS and $d$ is the cultural index denoting different answers to the same question $x$. For example, for a question $x$=“Do you agree with on the whole, men make better political leaders than women do?”, the answer $y=\text{Disagree}$ if $d=\text{English}$; and $y=\text{Strongly agree}$ if $d=\text{Arabic}$. Therefore, we only augment the question $x$ to be $x^{\prime}$ but retain the same opinion $y$ as the original $x$. We further denote vanilla LLM and CultureLLM as $f$ and $f^{\star}$, respectively. Then, denoting $\ell$ as the loss function, our learning objective is formulated as:
$f^{\star}=\arg\min_{f}\mathbb{E}_{\begin{subarray}{c}(x_{j},y_{j}^{d})\in\mathcal{D}_{d}\\ (x^{\prime}_{j},y_{j}^{d})\in g(\mathcal{D}_{d})\end{subarray}}[\ell(f(x_{j}),y_{j}^{d})+\ell(f(x^{\prime}_{j}),y_{j}^{d})].$
@@
3.2 Sampling
The sampling process should follow two principles: 1) cover as many cultural topics as possible and 2) sample questions that can be clearly answered by LLMs. Based on the two principles, we manually select $n=50$ questions and rewrite them into the Question-Answer (QA) format, covering $7$ topics, namely social values, security, science and technology, religious values, ethical values and norms, political interest and political participation, and migration.
@@ 3.3 Semantic Data Augmentation
The samples from WVS are not enough to fine-tune, which are then augmented by our semantic data augmentation approach. In a formal sense, semantic augmentation retains the original ground-truth opinions ($y_{d}$) from different cultures and only generates semantically equivalent questions ($x$). A naive augmentation approach is to directly use strong LLMs such as GPT-4 to generate new samples (Walters and Wilder, 2023), which could introduce mode collapse, as generation quality can only be controlled by prompts. Furthermore, since LLMs could suffer from cultural bias, directly generating cultural data using prompts could lead to unexpected or even erroneous outputs.
@@
As shown in Figure 2, the augmentation consists of two stages: semantic template generation and intact sentence generation. The first stage generates several semantically equivalent but stylistically different sentences and parses them into semantic templates. The second stage then generates samples by replacing certain words in the semantic templates. Such an augmentation introduces more diversities: The first stage introduces sentence-level diversities, and the second one introduces word-level diversities.
3.3.1 Semantic Template Generation
@@
This stage generates semantically equivalent question templates $\mathcal{T}=\{t_{i}\}_{i=1}^{k}$ based on $x\in\mathcal{D}_{d}$. The generation process is nontrivial since there are two challenges ahead: 1) the naturalness and diversities and 2) the semantic preservation. We solve the first challenge by leveraging GPT-4 as the generator with certain prompts to ensure naturalness and diversity. Then, we solve the second challenge by introducing a semantic preservation filter $\mathcal{P}$ to measure the similarity between the original and generated sentences. [^7] For computational efficiency, we use BERT embedding as $\mathcal{P}$ while other models can also be used.
Specifically, we first use the prompt “Could you please generate [$n$] sentences that (1) have different sentence structures and (2) have the same meaning with the following sentence: $x_{i}$” to generate $n$ sentences using GPT-4. Then, we denote the embedding of the original sentence and the generated sentences as $z=\mathcal{P}(x)$ and $z^{\prime}=\mathcal{P}(t)$, respectively. Then we compute their similarity score $c=\cos(z,z^{,})$. If $c$ passes the threshold value $\tau$, the generated sentence will be reserved:
$\mathcal{T}=\{t_{i}|\cos(\mathcal{P}(t_{i}),\mathcal{P}(x_{j}))>\tau\},\forall x_{j}\in\mathcal{D}_{d}.$
@@
Specifically, for sample “Do you agree with One of my main goals in life has been to make my parents proud?”, we generate $m$ samples using GPT-4, which are then go through the semantic filter $\mathcal{P}$ to eventually retain $k(k\leq m)$ semantically equivalent sentences, e.g. “Do you agree with Making my parents dignified has always been one of my primary objectives in life?" and “Do you agree with The goal to bring vanity to my parents has been a central life goal of mine?"
@@
To make the generated data more diverse, we then parse the $n$ sentences to find the proper components to replace, which construct the templates. For efficiency and cost-saving, we use NLTK (Loper and Bird, 2002) to find replaceable words, such as adjectives, adverbs, nouns and verbs. The semantic templates are like “Do you agree with The [x] to bring [x] to my parents has been a [x] life [x] of mine?" where “[x]” is the replaceable part. In total, we generate $k$ templates for each sample $x_{j}\in\mathcal{D}_{d}$.
3.3.2 Intact Sample Generation
@@
This step is to randomly replace synonyms in templates to generate fine-tuning samples. Synonyms are diverse in different contexts, and existing synonym tables could ignore their contexts, thus we apply GPT-4 to generate context-aware synonyms for words in the templates and randomly replace some of them. To further preserve semantics, we also use the semantic preservation filter in this step. After filtering, we generate $m$ samples for each template $t_{i}\in\mathcal{T}$, and get $n^{\prime}=mnk$ samples for all $x_{j}\in\mathcal{D}_{d}$ in total.
For instance, the intact samples for template “Do you agree with The [x] to bring [x] to my parents has been a [x] life [x] of mine?" could be “Do you agree with The goal to bring pride to my parents has been a central life goal of mine?" and “Do you agree with The hope to bring vanity to my parents has been a central life goal of mine?"
Our human study in Section 4.6 shows that semantic data augmentation can generate high-quality and semantically equivalent sentences.
3.4 Converting WVS opinions to groundtruth

### v2

@@ Abstract
Large language models (LLMs) have been observed to exhibit bias towards certain cultures due to the predominance of training data obtained from English corpora. Considering that multilingual cultural data is often expensive to procure, existing methodologies address this challenge through prompt engineering or culture-specific pre-training. However, these strategies may neglect the knowledge deficiency of low-resource cultures and necessitate substantial computing resources. In this paper, we propose CultureLLM, a cost-effective solution to integrate cultural differences into LLMs. CultureLLM employs the World Value Survey (WVS) as seed data and generates semantically equivalent training data through the proposed semantic data augmentation. Utilizing only $50$ seed samples from WVS with augmented data, we fine-tune culture-specific LLMs as well as a unified model (CultureLLM-One) for $9$ cultures, encompassing both rich and low-resource languages. Extensive experiments conducted on $60$ culture-related datasets reveal that CultureLLM significantly surpasses various counterparts such as GPT-3.5 (by $8.1$%) and Gemini Pro (by $9.5$%), demonstrating performance comparable to or exceeding that of GPT-4. Our human study indicates that the generated samples maintain semantic equivalence to the original samples, offering an effective solution for LLMs augmentation. Code is released at https://github.com/Scarelette/CultureLLM.
@@ 1 Introduction
Culture is a complex construct that encapsulates various identities, including, but not limited to, language, nationality, region, religion, and gender identity. Cultural bias is prevalent worldwide and refers to the tendency to favor specific cultural perspectives, values, and norms, which results in subjective opinions and may offend individuals from other cultures. For instance, according to the World Value Survey (Survey, 2022), Arabic culture believes that men are better political leaders than women, while people in the United States maintain a contrary opinion. [^1] We respect all opinions in different cultures. As large language models (LLMs) (OpenAI, 2023b; Google, 2023) gain prominence, they are reported to exhibit cultural bias and specifically show partiality towards Western culture, as English corpora dominate the training data (Liu et al., 2023b; Cao et al., 2023; Masoud et al., 2023; Naous et al., 2023; Wang et al., 2023d; Johnson et al., 2022). Low-resource cultures are frequently underrepresented due to the insufficient training data available from these cultures. LLMs’ cultural bias constitutes a significant bottleneck in human-AI collaboration and considerably impedes AI democracy.
@@
Tackling cultural bias necessitates that a large language model (LLM) acknowledges cultural differences (Hofstede, 1984). Kovavc et al. Kovač et al. (2023) and Want et al. Wang et al. (2023d) thought LLMs have enough knowledge of all cultures and devised prompt engineering technologies to induce LLMs to exhibit specific cultural perspectives. However, they are not effective, especially in low-resource cultures with limited data. Another line of work pre-trained culturally aware LLMs and then fine-tuned on specific datasets (Chan et al., 2023; Nguyen et al., 2023b; Pipatanakul et al., 2023; Abbasi et al., 2023; Lin and Chen, 2023). They require the collection of large-scale pre-training and fine-tuning datasets and extensive computing resources, thus are not affordable to ordinary researchers and cannot handle low-resource culture. To date, training culturally aware LLMs at affordable costs remains a challenge.
Figure 1: Overview of CultureLLM. CultureLLM consists of three steps: sampling, semantic data augmentation, and fine-tuning. Culture-specific and unified CultureLLM can be fine-tuned.
@@
In this paper, we propose CultureLLM, a cost-effective [^2] Fine-tuning a CultureLLM only costs \$6 via OpenAI API. solution to incorporate cultural differences into LLMs. Technically speaking, CultureLLM is inspired by the well-known fact that LLMs are inevitably not robust to the style and format of the prompts (Zhu et al., 2023), indicating that we can further leverage such a weakness to further improve the performance of LLMs by enriching the prompts. In particular, we focus on cultural values in this work. As shown in Figure 1, CultureLLM consists of three steps: sampling, semantic data augmentation, and fine-tuning. Inspired by Attitude-Behavior Consistency theory (Fazio and Zanna, 1981) which emphasizes that people’s opinion is consistent with their behaviors, we use the World Values Survey (WVS) (Survey, 2022) as seed data. Then, we devise a semantic data augmentation approach to generate semantically equivalent samples. The aim is to generate semantic equivalent inputs, thus we can get the ground-truth from seed data directly. Finally, CultureLLM is obtained by fine-tuning both the seed and the generated data. WVS is a public opinion poll that contains people’s opinions on cultural topics from different countries. To be specific, we select $50$ seed samples from WVS, covering $7$ topics: “social values", “migration", “security", “science and technology", “religious values", “ethical values and norms", and “political interest and political participation". Using these generated samples and answers from people in different cultures, we fine-tune specific and unified LLMs: specific LLMs are tailored for each culture such as CultureLLM-Ar for Arabic and CultureLLM-Tr for Turkish; unified LLMs (CultureLLM-One) are one LLM that fits all cultures. [^3] The bound of culture is unclear; we use the main spoken language to distinguish cultures (Delanoy, 2020b).
@@
We build $9$ specific CultureLLM and a CultureLLM-One covering both high- and low-resource cultures: Arabic culture, Bengali culture, Chinese culture, English culture, German culture, Korean culture, Portuguese culture, Spanish culture, and Turkish culture. Then we evaluated them on $8$ culture-related downstream tasks: offensive language detection, hate speech detection, stance detection, toxicity detection, threat detection, bias detection, abusive detection, spam detection, and an open-ended generative task. We have $60$ test sets of $68,672$ samples in total. Experiments show that CultureLLM fine-tuned on GPT-3.5 significantly outperforms GPT-3.5 by $8.1$% and outperforms Gemini pro (Google, 2023) by $9.5$% on average F1 score, achieving comparable or even better performance with GPT-4. Our human study of $50$ people demonstrates that the augmentation method can generate semantically equivalent samples. We further interpret the rationale behind its effectiveness by exploring the fine-tuning data size and case studies. Finally, results on Big-Bench Hard (Suzgun et al., 2022) and GSM8K (Cobbe et al., 2021) indicate that CultureLLM is resistant to catastrophic forgetting. CultureLLM also supports fine-tuning LLMs of open-source models.
@@
    We conducted extensive experiments in a wide range of cultures and LLMs, showing that LLMs performs consistently well in all downstream tasks.
@@ 2.1 Cultural Problem and Solution in LLMs
The ideal solution is to improve the cultural awareness of LLMs. There are mainly two types of approach: prompt engineering and pre-training. Kovač et al. (2023); Wang et al. (2023d) thought LLMs as superpositions of cultural perspectives, which can be prompted to targeted cultural perspectives. while Rao et al. (2023) encoded cultural values in the prompts. Although PE is cheap, its effectiveness is challenged, especially in low-resource cultures where LLMs lack such cultural knowledge due to lack of representation in pre-training data. Another line of research is pre-training and fine-tuning (Chan et al., 2023; Nguyen et al., 2023b; Pipatanakul et al., 2023; Abbasi et al., 2023; Lin and Chen, 2023) that trains culturally-aware LLMs for different cultures by collecting large-scale pre-training datasets and then performed fine-tuning for better alignment. While they achieved great performance, this approach is too expensive and time-consuming, thus it is difficult to apply to more cultures and countries. They still suffer from a low-resource culture problem where the pre-training data are difficult to collect. MaLA-500 (Lin et al., 2024) trained a new LLM on Llama 2 to cover $534$ languages, which is resource intensive.
@@ 2.2 Data Augmentation for LLMs
Human-annotated data are high-quality but expensive. Due to the strong generation ability of LLMs, many works focused on leveraging data augmentation for LLMs. Yu et al. (2023); Liu et al. (2023a) used LLMs to augment the math data and then fine-tuned with those data. Li et al. (2023) synthesized data with two designed modules: self-augmentation and self-curation. Chen et al. (2024) introduced a self-play mechanism, where LLM generates its own training data from its previous iterations, refining its policy by discerning these self-generated responses from those obtained from human-annotated data. There are also other uses for synthetic data, such as knowledge distillation (Wang et al., 2023c) and improving text embedding tasks (Wang et al., 2023a). Our data augmentation approach also adopts LLMs for data generation, but we add controllable modules such as template editing, synonym replacement, and semantic filter to ensure the diversity and semantic equivalence of the generated samples. It can also be used as a general augmentation method in other applications.
Efforts in cultural datasets (Nguyen et al., 2023a; Fung et al., 2022) focus on cultural common sense and norms. However, they generate data from only the English or Chinese corpus and thus may contain cultural bias toward other cultures. In contrast, World Values Survey (WVS) (Survey, 2022) is a large-scale pool that contains answers from people of different cultures, thus providing more objective cultural values from specific cultures.
This work is also related to value alignment (Ji et al., 2023; Shen et al., 2023; Yao et al., 2023) to align the values of LLMs with human’s by designing algorithms for value measurement and behavior alignment. In contrast, this work primarily emphasizes value understanding with the potential to be extended for value alignment. For instance, semantic augmentation can be used to generate training data for alignment-related tasks.
@@ 3.1 Overview
Cultural differences are prevalent in various cultures and backgrounds, leading to an impact on outcomes in downstream applications such as hate speech and biased language. To address the gap between low-source cultural data collection and its wide applications, we design CultureLLM by fine-tuning an LLM on data generated by our novel semantic data augmentation approach leveraging the sensitivity of LLMs on prompts (Zhu et al., 2023). Figure 1 presents an overview of CultureLLM, where the first step is to sample a subset of data from an existing World Value Survey (WVS) (Survey, 2022) that represents different opinions (answers) towards the same value questions given by native users. The adoption of WVS is inspired by Attitude-Behavior Consistency theory (Fazio and Zanna, 1981), which emphasizes the strong relationship between attitude and behavior. Therefore, WVS serves as an ideal seed for data augmentation. [^4] WVS is one feasible option for seed and other surveys can also be used. But the cultural survey data are extremely rare and the WVS could be the most comprehensive one. More details about WVS are in Section B.1. After sampling, the second step is to generate augmented data using our proposed semantic augmentation approach (Section 3.3) and then fine-tune a CultureLLM for each specific culture such as CultureLLM-Ar for Arabic culture and CultureLLM-Tr for Turkish culture.
Generally speaking, we use $\mathcal{D}_{d}=\{(x_{j},y_{j}^{d})\}_{j=1}^{n}$ to denote the seed and $\mathcal{D}^{\prime}_{d}=g(\mathcal{D}_{d})=\{(x^{\prime}_{j},y_{j}^{d})\}_{j=1}^{n^{\prime}}$ as the augmented data with $g(\cdot)$ the augmentation algorithm. Note that the question $x$ here is the same in all cultures in WVS and $d$ is the cultural index denoting different answers to the same question $x$. For example, for a question $x$=“Do you agree with on the whole, men make better political leaders than women do?”, the answer $y=\text{{\footnotesize Disagree}}$ if $d=\text{{\footnotesize English}}$; and $y=\text{{\footnotesize Strongly agree}}$ if $d=\text{{\footnotesize Arabic}}$. Therefore, we only augment the question $x$ to be $x^{\prime}$ but retain the same opinion $y$ as the original $x$. We also denote vanilla LLM and CultureLLM as $f$ and $f^{\star}$, respectively. Then, denoting $\ell$ as the loss function, our learning objective is formulated as: $f^{\star}=\arg\min_{f}\mathbb{E}_{\begin{subarray}{c}(x_{j},y_{j}^{d})\in\{\mathcal{D}_{d},g(\mathcal{D}_{d})\}\end{subarray}}[\ell(f(x_{j}),y_{j}^{d})].$
3.2 Sampling
@@
The sampling process should follow two principles: 1) cover as many cultural topics as possible and 2) sample questions that can be clearly answered by LLMs. Based on the two principles, we manually select $n=50$ questions and rewrite them in the Question-Answer (QA) format, covering $7$ topics, namely social values, security, science and technology, religious values, ethical values and norms, political interest and political participation, and migration. The details can be found in Section B.1.
@@ 3.3 Semantic Data Augmentation
Samples from WVS are not enough to fine-tune, which can be augmented by our semantic augmentation approach. In a formal sense, semantic augmentation retains the original ground-truth opinions ($y_{d}$) from different cultures and only generates semantically equivalent questions ($x$). A naive augmentation approach is to directly use strong LLMs such as GPT-4 to generate new samples (Walters and Wilder, 2023), which could introduce mode collapse, as generation quality can only be controlled by prompts. Furthermore, since LLMs could suffer from cultural bias, directly generating cultural data using prompts could lead to unexpected or even erroneous results.
@@
As shown in Figure 2, the augmentation consists of two stages: semantic template generation and intact sentence generation. The first stage generates several semantically equivalent but stylistically different sentences and parses them into semantic templates. The second stage then generates samples by replacing certain words in the semantic templates. Such an augmentation can naturally introduce more diversities: The first stage increases sentence-level diversities and the second improves the word-level diversities.
Semantic Template Generation
@@
This stage generates semantically equivalent question templates $\mathcal{T}=\{t_{i}\}_{i=1}^{k}$ based on $x\in\mathcal{D}_{d}$. The generation process is nontrivial since there are two challenges ahead: 1) the naturalness and diversities and 2) the semantic preservation. We solve the first challenge by using GPT-4 as the generator with certain prompts to ensure naturalness and diversity. Then, we solve the second challenge by introducing a semantic preservation filter $\mathcal{P}$ to measure the similarity between the original and generated sentences. [^5] For computational efficiency, we use BERT embedding as $\mathcal{P}$ while other models can also be used.
@@
We first use the prompt “Could you please generate [$n$] sentences that (1) have different sentence structures and (2) have the same meaning with the following sentence: $x_{i}$” to generate $n$ sentences using GPT-4. Then, we denote the embedding of the original sentence and the generated sentences as $z=\mathcal{P}(x)$ and $z^{\prime}=\mathcal{P}(t)$, respectively. Then we compute their similarity score $c=\cos(z,z^{,})$. If $c$ passes the threshold value $\tau$, the generated sentence will be reserved:$\mathcal{T}=\{t_{i}|\cos(\mathcal{P}(t_{i}),\mathcal{P}(x_{j}))>\tau\},\forall x_{j}\in\mathcal{D}_{d}.$ Specifically, for sample “Do you agree with One of my main goals in life has been to make my parents proud?”, we generate $m$ samples using GPT-4, which are then go through the semantic filter $\mathcal{P}$ to eventually retain $k(k\leq m)$ semantically equivalent sentences, e.g. “Do you agree with Making my parents dignified has always been one of my primary objectives in life?" and “Do you agree with The goal to bring vanity to my parents has been a central life goal of mine?"
@@
To diversify the generated data, we parse the $n$ sentences to find the appropriate components to replace, which construct the templates. For efficiency, we use NLTK (Loper and Bird, 2002) to find replaceable words, such as adjectives, adverbs, nouns, and verbs. The semantic templates are like “Do you agree with The [x] to bring [x] to my parents has been a [x] life [x] of mine?" where “[x]” is the replaceable part. In total, we generate $k$ templates for each sample $x_{j}\in\mathcal{D}_{d}$.
Intact Sample Generation
@@
This step is to replace synonyms in templates to generate fine-tuning samples. We apply GPT-4 to generate context-aware synonyms in the templates and randomly replace some of them. To further preserve semantics, we also use the semantic preservation filter. After filtering, we generate $m$ samples for each template $t_{i}\in\mathcal{T}$, and get $n^{\prime}=mnk$ samples for all $x_{j}\in\mathcal{D}_{d}$ in total. For example, intact samples for template “Do you agree with The [x] to bring [x] to my parents has been a [x] life [x] of mine?" could be “Do you agree with The goal to bring pride to my parents has been a central life goal of mine?" and “Do you agree with The hope to bring vanity to my parents has been a central life goal of mine?" Our human study in Section 4.6 shows that augmentation can generate high-quality and semantically equivalent sentences.
3.4 Fine-tuning